f 150cm3 of steam at 100∘C is added to 250cm3 water at 40∘C, how much steam(in cm3) is condensed into water at 100∘C?
1
Expert's answer
2020-12-10T11:09:30-0500

"V_1=150 ~sm^3,~V_2=250 ~sm^3,~t_1=100\u00b0C,~t_2=40\u00b0C,"

"L=2.3\cdot10^6~\frac{J}{kg},~c=4200~\frac{J}{kg\cdot °C}, \rho_1=0.6~\frac{kg}{m^3},~\rho_2=1000~\frac{kg}{m^3}."


"Q_1=L\rho_1V_1=207~J."

"Q_2=c\rho_2V_2(t_1-t_2)=63000~J."

"Q_1\ll Q_2," so full steam ("150 ~sm^3") is condensed into water at 100°C.